OpenStudy (anonymous):

Which of the following circles have their centers on the x-axis? Check all that apply. A. (x - 0)2 + (y - 5)2 = 49 B. (x - 0)2 + (y - 0)2 = 25 C. (x - 4)2 + (y - 0)2 = 22 D. (x + 1)2 + (y - 7)2 = 16

OpenStudy (anonymous):

@ganeshie8

OpenStudy (xapproachesinfinity):

apply the same thing here as well

OpenStudy (anonymous):

You would want to find where the y value of the equations is equal to zero. A little trick I do is like this. In your letters A, and D, we see that the y function is (y-5) and (y-7) respectively. What that is saying is that the y coordinates of those two are 5 and 7, because 5-5 = 0 and 7-7=0. So, which of the y functions will be equal to zero? For B and C we see that they are (y-0) and (y-0), if y = 0, then we must be on the x axis. Therefore, B and C are functions of the circles located on the x axis.